I spoke with the teacher - they said the "see notes" isnt needed

Columbia Sportswear Company is a leader in outdoor sportswear. The following are Columbias financial statements as presentedCOLUMBIA SPORTSWEAR COMPANY CONSOLIDATED STATEMENTS OF COMPREHENSIVE INCOME (In thousands) 2016 Year Ended December 31. 2015COLUMBIA SPORTSWEAR COMPANY CONSOLIDATED BALANCE SHEETS (In thousands) December 31, 2016 2015 ASSETS UNIWINU EIEEEEEEEEEEEEEECONSOLIDATED STATEMENTS OF CASH FLOWS (In thousands) Year Ended December 31. 2015 2016 2014 198,439 $ 179,851 $ 141,859 60,01COLUMBIA SPORTSWEAR COMPANY CONSOLIDATED STATEMENTS OF EQUITY (In thousands) Columbia Sportswear Company Shareholders EquityExpand Your Critical Thinking 12-02 a Columbia Sportswear Companys financial statements are presented in Appendix B. Click h

VF CORPORATION Consolidated Balance Sheets December 2016 I thousands exeplaremont 1.227.862 $ 944.423 1,197,678 1.569.325 298VF CORPORATION Consolidated Statements of Income 2014 S Ishands, 11.900,314 S 116.699 12.019,0X13 per 11.09.15 $ 123,020 12.0See notes to consolidated financial statements VF CORPORATION Consolidated Statements of Comprehensive Income Year Ended Dece2016 2015 2014 In thousands Operating activities Net income $ 1,074,106 $ 1,231,593 S 1,047,505 Adjustments to reconcile netVF CORPORATION Consolidated Statements of Stockholders Equity Accumulated Additional Other Common Stock Paldin Comprehensive
Answer #1

Free Cash flow = cash flow from operating activities - capital expenditures

Columbia sportswear company = 275167-49987 = $255180

VF Corporation = 1477919-175840=$1302079